When Katie was born her mother invested $5000 in an account for her college savings. The interest rate is 3.5% compounded annually. To represent this, we can use the formula V = 5000(1+r)t where r represents the interest rate and t represents the time in years. How much will Katie have in her account when she turns 18?
A. 5,175
B. 7,127
C. 9,287
D. 12,472

Answer by stanbon(75887)

When Katie was born her mother invested $5000 in an account for her college savings. The interest rate is 3.5% compounded annually.
To represent this, we can use the formula V = 5000(1+r)t where r represents the interest rate and t represents the time in years. How much will Katie have in her account when she turns 18?
---
V(18) = 5000(1.035)^18 = $9287
===================================
Cheers,
Stan H.
==================
A. 5,175
B. 7,127
C. 9,287
D. 12,472